# Detecting Knowledge Boundary of Vision Large Language Models by Sampling-Based Inference

### v1

@@ Abstract
Despite the advancements made in Visual Large Language Models (VLLMs), like text Large Language Models (LLMs), they have limitations in addressing questions that require real-time information or are knowledge-intensive. Indiscriminately adopting Retrieval Augmented Generation (RAG) techniques is an effective yet expensive way to enable models to answer queries beyond their knowledge scopes. To mitigate the dependence on retrieval and simultaneously maintain, or even improve, the performance benefits provided by retrieval, we propose a method to detect the knowledge boundary of VLLMs, allowing for more efficient use of techniques like RAG. Specifically, we propose a method with two variants that fine-tunes a VLLM on an automatically constructed dataset for boundary identification. Experimental results on various types of Visual Question Answering datasets show that our method successfully depicts a VLLM’s knowledge boundary based on which we are able to reduce indiscriminate retrieval while maintaining or improving the performance. In addition, we show that the knowledge boundary identified by our method for one VLLM can be used as a surrogate boundary for other VLLMs. Code will be released at https://github.com/Chord-Chen-30/VLLM-KnowledgeBoundary
@@ 1 Introduction
The great advancements in language models have led to the integration of image encoding and understanding capabilities Achiam et al. (2023); Lu et al. (2024); Wang et al. (2024), significantly enhancing the performance of a series of pre-trained Visual Large Language Models (VLLMs) in tasks involving Visual Question Answering (VQA). Despite these advancements, akin to text Large Language Models (LLMs) Touvron et al. (2023); Workshop et al. (2022); Zhang et al. (2024b), VLLMs remain constrained by the boundaries of their knowledge Lin and Byrne (2022). As a result, their ability to accurately respond to content outside the model’s knowledge scope, such as knowledge-intensive questions, real-time news, and queries with dynamic answers, is considerably limited.
@@ 2.3 Training
Setting a threshold to binarily classify the queries might be an overly rigid method and there is no room for adjustment when the knowledge boundary model performs poorly in possibly unseen scenarios unless we adjust $\epsilon$ and retrain the model. Thus, we also propose a method that can depict a softer boundary. Recall that for query $\bm{q}$, the average score $s$ over $R$ model predictions ranges in $[s_{w},s_{c}]$, where $s_{w}$ indicates a wrong answer and $s_{c}$ indicates a correct one. We linearly flip the score, for example, the new score $s^{{}^{\prime}}=s_{w}$ represents a strong tendency for external knowledge while $s^{{}^{\prime}}=s_{c}$ represents a refusal to external knowledge.
@@ 3.1.2 Test Data
Table 2: Test data property illustration that whether RAG is helpful in answering the queries.
@@ 3.1.3 Use of RAG
We aim not only to locate the queries that need RAG to answer better but also adopt retrieval techniques to verify the final VQA performance with the search decision $HKB_{\phi}$ and $SKB_{\phi}$ defined in equations 6. We note that although there are various options for retrieval, such as text search and image search, we do not design detailed methods to determine the best option in this paper. Instead, we directly use text search (Google) for Dyn-VQA and image search (Bing) for the rest for better retrieval information quality towards answering the question. We note that Dyn-VQA is a challenging dataset that includes many samples of multi-hop property, therefore we use the golden query Li et al. (2024) have summarized for retrieving useful information.
@@ 3.1.4 Base Models
For Visual Question Answering, we first evaluate the performance of the original models to be sampled. In addition, we seek to validate whether the identified knowledge boundary can function as a surrogate boundary for other VLLMs since constructing training datasets through sampling (Sec. 2.3) on (larger) models can be prohibitively expensive. We further validate the surrogate knowledge boundary on the following VLLMs, Qwen-VL-Max Bai et al. (2023), Qwen-VL-2 Wang et al. (2024) and GPT4-o Hurst et al. (2024), to evaluate its potential for generalizing across different VLLMs.
@@ 3.2 Main Results
We present our main results of Qwen-VL-7B-Chat in Table 3 and result of DeepSeek-VL-7B-Chat in Appendix A.4. In this section, we focus on the results of Qwen.
Metrics LLM represents that the score is evaluated by a text LLM, Qwen-Max, given the model prediction and gold answer. Metrics Acc. refers to token accuracy which involves determining the proportion of tokens in the model’s predictions that match the tokens in the gold answer. Both Scores range from 0 to 100 and a higher score indicates a higher performance. The % columns refer to the ratio of data that our knowledge boundary model predicts to lie beyond the VLLM’s knowledge boundaries. The “Human” column represents the corresponding statistics where the Knowledge Boundary model is trained on the human-labeled data mentioned in Sec. 3.1.1 and we deem it a reference result.
@@ 4 Analysis
In this section, we present three analytical experiments. The first one shows the performance of other VLLMs if we employ the identified knowledge boundary as a surrogate. The second shows how the RAG ratio and VQA performance are affected by the threshold defined in the $SKB$ variant. The third one presents the accuracy of VLLM boundary identification on held-in data at training time.
@@ 4.1 Surrogate Boundary for Other VLLMs
We assemble around 340 thousand VQA samples from various domains discussed in Sec. 3.1.1. Sampling each data thirty times is prohibitively expensive for closed-source VLLMs. Although different VLLMs intuitively possess varying scopes of knowledge, we believe that it is highly probable that a significant portion of these scopes overlap. For instance, queries regarding recently occurring news events typically fall outside the knowledge boundaries of any model. Thus, we conduct an experiment that validates whether the identified knowledge boundary can function as a surrogate boundary for other VLLMs.
@@
From Table 4 Mix row, Qwen-VL-Max, Qwen-VL-2 and GPT4-o achieve better performance than all three baseline settings. Deepseek-VL-7B-Chat remains competitive to the “All RAG” setting with LLM metric and outperforms all other settings in Table 9 Mix row. For other datasets, we show that the previously identified knowledge boundary can help maintain the performance with a reduced RAG ratio. For example, GPT4-o achieves 54.83 with only 61.74% RAG ratio while the “All RAG” setting achieves 55.47 on the Life VQA dataset. Deepseek-VL-7B-Chat maintains its performance on the Dyn-VQA (en) dataset compared to the “All RAG” setting and keeps a clear margin compared to the “No RAG” setting with a 23.92% retrieving deduction.
4.2 Effect of $\epsilon$ for $SKB$
@@ 4.3 Efficiency
Our method incorporates an additional forward pass for each VQA example for knowledge boundary identification. We report the overall efficiency in Table 5 on the Mix dataset, where the All RAG setting always uses RAG (calls to search engine included) and does not perform the forward pass, and $HKB/SKB$ refers to partially performing RAG according to our model’s predictions with forward-pass time included.
@@ 4.4 Performance of Knowledge Boundary Identification on Held-In Data
Table 6: Training and validation results on the held-in dataset. Metrics are shown in the accuracy defined in swift package. We have a limited number of human-labeled samples thus we do not set a validation set for “Human-labeled” setting.
@@ 6 Conclusion
In this paper, we introduce a method with two variants that fine-tunes VLLMs on automatically constructed datasets for boundary identification. This method mitigates the reliance on RAG techniques, which introduce significant latency and long input sequences. Our experiments across diverse held-out VQA datasets, including knowledge-intensive, non-knowledge-intensive, and mixed datasets, demonstrate that our method not only maintains or enhances VLLM performance but also lowers the RAG ratio. Additionally, the fine-tuned knowledge boundary exhibits versatility by functioning as a surrogate for other VLLM series, facilitating retrieval reduction without compromising performance. These findings underscore the efficacy of our approach in optimizing the balance between retrieval dependence and model performance, paving the way for more efficient and effective deployment of VLLMs in practical applications.
@@ 7 Limitations
In this paper, we do not design detailed methods to distinguish the search type, such as text search and image search, towards answering a VQA sample. Experiments utilizing training data sampled from larger VLLMs are currently lacking. Both limitations will be addressed in our future work.
@@ A.3 Training Details and Hyperparameters
Table 7: Detailed hyperparameters.
@@
A.4 Main Results on DeepSeek
@@
A.5 Supplementary Results of “Surrogate Boundary” Experiments
We provide the supplementary experimental results for Sec 4.1 where the token accuracy metrics are shown in Table 9. It can be concluded that similar conclusions can be drawn as in Sec 4.1. The experiment where DeepSeek-VL-7B-Chat is trained for surrogate boundary prediction is shown in Table 10 and 11.
@@ A.5 Supplementary Results of “Surrogate Boundary” Experiments
GPT4-o 34.52 35.96 35.17 13.50% 35.77 25.00% 35.86 54.83% 36.24 74.50%
Table 11: Knowledge Boundary model (DeepSeek-VL-7B-Chat) as a surrogate boundary identifier for other VLLMs. Results evaluated by token accuracy.
A.6 Supplementary Results on MMMU Dataset
@@
The results in Table 12 show the Knowledge Boundary model trained by human-labeled data helps achieve the best performance. It verifies that the aforementioned Human-labeled training data is effective. In addition, we show that our methods also exhibit substantial potential within this setting, in which both the $HKB$ and $SKB$ models predict a high search ratio over MMMU. We contend that the suboptimal performance of this dataset arises because it lies beyond the knowledge boundaries, that are challenging to validate using RAG, as delineated by the white dashed lines in Fig. 1. We present the performance of each of the 30 subjects in the MMMU validation set in Fig 4. The first row shows the LLM evaluation results and the second shows the token accuracy metric. We can see that in most subjects “Human” setting succeeds in obtaining a higher performance than both “All RAG” and “No RAG” settings.
@@
Figure 4: Qwen-VL-Max and Qwen-VL-2 performance on MMMU validation set with Knowledge Boundary model trained on Human-labeled data.

### v3

@@ Abstract
Despite the advancements made in Vision Large Language Models (VLLMs), like text Large Language Models (LLMs), they have limitations in addressing questions that require real-time information or are knowledge-intensive. Indiscriminately adopting Retrieval Augmented Generation (RAG) techniques is an effective yet expensive way to enable models to answer queries beyond their knowledge scopes. To mitigate the dependence on retrieval and simultaneously maintain, or even improve, the performance benefits provided by retrieval, we propose a method to detect the knowledge boundary of VLLMs, allowing for more efficient use of techniques like RAG. Specifically, we propose a method with two variants that fine-tune a VLLM on an automatically constructed dataset for boundary identification. Experimental results on various types of Visual Question Answering datasets show that our method successfully depicts a VLLM’s knowledge boundary, based on which we are able to reduce indiscriminate retrieval while maintaining or improving the performance. In addition, we show that the knowledge boundary identified by our method for one VLLM can be used as a surrogate boundary for other VLLMs. Code will be released at https://github.com/Chord-Chen-30/VLLM-KnowledgeBoundary
@@ 1 Introduction
The great advancements in language models have led to the integration of image encoding and understanding capabilities Achiam et al. (2023); Lu et al. (2024); Wang et al. (2024), significantly enhancing the performance of a series of pre-trained Vision Large Language Models (VLLMs) in tasks involving Visual Question Answering (VQA). Despite these advancements, akin to Large Language Models (LLMs) Touvron et al. (2023); Workshop et al. (2022); Brown et al. (2020); Zhang et al. (2024b), VLLMs remain constrained by the boundaries of their knowledge Lin and Byrne (2022). As a result, their ability to accurately respond to content outside the model’s knowledge scope, such as knowledge-intensive questions, real-time news, and queries with dynamic answers, is considerably limited.
@@ 2.3 Training
Setting a threshold to binarily classify the queries might be an overly rigid method and there is no room for adjustment when the knowledge boundary model performs poorly in possibly unseen scenarios unless we adjust $\epsilon$ and retrain the model. Thus, we also propose a method that can depict a softer boundary. Recall that for query $\bm{q}$, the average score $s$ over $R$ model predictions ranges in $s_{w},s_{c}]$, where $s_{w}$ indicates a wrong answer and $s_{c}$ indicates a correct one. We linearly flip the score, for example, the new score $s^{{}^{\prime}}=s_{w}$ represents a strong tendency for external knowledge while $s^{{}^{\prime}}=s_{c}$ represents a refusal to external knowledge.
@@ 3.1.2 Test Data
Table 2: Test data property illustration of whether RAG is helpful in answering the queries.
@@ 3.1.3 Use of RAG
We aim not only to locate the queries that need RAG to answer better but also to adopt retrieval techniques to verify the final VQA performance with the search decision $HKB_{\phi}$ and $SKB_{\phi}$ defined in equations 6. We note that although there are various options for retrieval, such as text search and image search, we do not design detailed methods to determine the best option in this paper. Instead, we directly use text search (Google) for Dyn-VQA and image search (Bing) for the rest for better retrieval information quality towards answering the question. We note that Dyn-VQA is a challenging dataset that exhibits multi-hop property, therefore we use the golden query Li et al. (2024) have summarized for retrieving useful information.
@@ 3.1.4 Base Models
For Visual Question Answering, we first evaluate the performance of the original models to be sampled. In addition, we seek to validate whether the identified knowledge boundary can function as a surrogate boundary for other VLLMs since constructing training datasets through sampling (Sec. 2.3) on (larger) models can be prohibitively expensive. We further validate the surrogate knowledge boundary on the following VLLMs, Qwen-VL-Max Bai et al. (2023), Qwen-VL-2 Wang et al. (2024) and GPT-4o Hurst et al. (2024), to evaluate its potential for generalizing across different VLLMs.
@@ 3.2 Main Results
We present our main results of Qwen-VL-7B-Chat in Table 3 and results of DeepSeek-VL-7B-Chat in Appendix A.6. In this section, we focus on the results of Qwen.
Metrics LLM represents that the score is evaluated by a text LLM, Qwen-Max, given the model prediction and gold answer. Metrics Acc. refers to token accuracy which involves determining the proportion of tokens in the model’s predictions that match the tokens in the gold answer. Both Scores range from 0 to 100 and a higher score indicates a higher performance. The % columns refer to the ratio of data that our knowledge boundary model predicts to lie beyond the VLLM’s knowledge boundaries. The “Human” column represents the corresponding statistics where the Knowledge Boundary model is trained on the human-labeled data mentioned in Sec. 3.1.1, and we deem it a reference result.
@@ 4 Analysis
In this section, we present five analytical experiments. The first shows the performance of other VLLMs if we employ the identified knowledge boundary as a surrogate. The second shows how the RAG ratio and VQA performance are affected by the threshold defined in the $SKB$ variant. The third presents the efficiency of our method. The fourth is a case study showing cases with different Knowledge Boundary model predictions (inside or outside the knowledge boundary). The last shows the accuracy of VLLM boundary identification on held-in data at training time.
@@ 4.1 Surrogate Boundary for Other VLLMs
We assemble around 340 thousand VQA samples from various domains discussed in Sec. 3.1.1. Sampling each data thirty times is prohibitively expensive for closed-source VLLMs. While VLLMs (e.g., Qwen-VL, DeepSeek-VL) may intuitively exhibit different knowledge scopes, this overlap is expected given their shared pretraining corpora (e.g., LAION Schuhmann et al. (2022), COCO Lin et al. (2014)), similar visual encoder architectures (e.g., CLIP variants), and common textual knowledge from large-scale web data. Besides, queries regarding recently occurring news events typically fall outside the knowledge boundaries of any model. Thus, we conduct an experiment that validates whether the identified knowledge boundary can function as a surrogate boundary for other VLLMs.
@@
From Table 4 Mix row, Qwen-VL-Max, Qwen-VL-2 and GPT-4o achieve better performance than all three baseline settings. Deepseek-VL-7B-Chat remains competitive to the “All RAG” setting with LLM metric and outperforms all other settings in Table 11 Mix row. For other datasets, we show that the previously identified knowledge boundary can help maintain the performance with a reduced RAG ratio. For example, GPT-4o achieves 54.83 with only 61.74% RAG ratio while the “All RAG” setting achieves 55.47 on the Life VQA dataset. Deepseek-VL-7B-Chat maintains its performance on the Dyn-VQA (en) dataset compared to the “All RAG” setting and keeps a clear margin compared to the “No RAG” setting with a 23.92% retrieving deduction.
4.2 Effect of $\epsilon$ for SKBSKB
@@
Cases where the model predicts to be out of the knowledge boundary Cases where the model predicts to be in the knowledge boundary
Question: How many World Series titles has the team won? [Uncaptioned image] Prediction w/o retrieval: The Houston Astros have won one World Series title, in 2017. $\rightarrow$ (incorrect) Prediction w/ retrieval: The Houston Astros have won two World Series titles. $\rightarrow$ (correct) (Source: Dyn-VQA (en)) Question: Can you add a caption to the image using a phrase? For example: A little girl in a white jacket and sandals. [Uncaptioned image] Prediction w/o retrieval: A blue jay perched on a tree branch. $\rightarrow$ (correct) Prediction w/ retrieval: A blue and white bird perched on a tree branch. $\rightarrow$ (correct) (Source: NoCaps)
Question: Who is the current Prime Minister of this country? [Uncaptioned image] Prediction w/o retrieval: The current Prime Minister of the United Kingdom is Rishi Sunak. $\rightarrow$ (incorrect) Prediction w/ retrieval: The current Prime Minister of the United Kingdom is Keir Starmer. $\rightarrow$ (correct) (Source: Dyn-VQA (en)) Question: How is the boat staying ashore? [Uncaptioned image] Prediction w/o retrieval: The boat is staying ashore by being tied to a wooden post with a yellow rope. The rope is wrapped around the post and secured, preventing the boat from drifting back into the water. $\rightarrow$ (correct) Prediction w/ retrieval: It is tied to a tree with rope. $\rightarrow$ (correct) (Source: Visual7W)
Table 5: Cases when the model predicts whether the VQA queries are in VLLMs’ knowledge boundary. Base model is Qwen-VL-Max.
@@ 4.3 Efficiency
Our method incorporates an additional forward pass for each VQA example for knowledge boundary identification. We report the overall efficiency in Table 6 on the Mix dataset, where the All RAG setting always uses RAG (calls to the search engines included) and does not perform the forward pass, and $HKB/SKB$ refers to partially performing RAG according to our model’s predictions with forward-pass time included.
4.4 Case Study
This section shows four cases with different Knowledge Boundary model predictions (inside or outside the knowledge boundary). In the left column of Table 5, the Knowledge Boundary model predicts that they are outside the knowledge boundary, and the retrieval indeed helps the model correct its response. In the right column, we show two example cases that our Knowledge Boundary model predicts to be in the knowledge boundary and thus do not need retrieval.
@@ 4.5 Performance of Knowledge Boundary Identification on Held-In Data
Table 7: Training and validation results on the held-in dataset. Metrics are shown in the accuracy defined in ms-swift package. We have a limited number of human-labeled samples thus we do not set a validation set for “Human-labeled” setting.
@@ 6 Conclusion
In this paper, we introduce a method with two variants that fine-tunes VLLMs on automatically constructed datasets for boundary identification. This method mitigates the reliance on RAG techniques, which introduce significant latency and long input sequences. Our experiments across diverse held-out VQA datasets, including knowledge-intensive, non-knowledge-intensive, and mixed datasets, demonstrate that our method not only maintains or enhances VLLM performance but also lowers the RAG ratio. Additionally, the fine-tuned knowledge boundary exhibits versatility by functioning as a surrogate for other VLLM series, facilitating retrieval reduction without compromising performance. These findings underscore the efficacy of our method in optimizing the balance between retrieval dependence and model performance, paving the way for more efficient and effective deployment of VLLMs in practical applications.
@@ 7 Limitations
In this paper, we do not design detailed methods to distinguish the search type, such as text search and image search, towards answering a VQA sample. Experiments utilizing training data sampled from larger VLLMs are currently lacking. Both limitations will be addressed in our future work. Moreover, our current method is not yet able to reliably distinguish errors stemming from visual misrecognition (e.g., incorrect object identification or multi-view ambiguity) from knowledge gaps.
Acknowledgment
This work was supported by Alibaba Group through Alibaba Innovative Research Program and the Core Facility Platform of Computer Science and Communication, SIST, ShanghaiTech University.
@@ A.3 Training Details and Hyperparameters
Table 8: Detailed hyperparameters. Grad. Accum. stands for gradient accumulation steps.
@@
A.4 Training and Evaluation Cost
Based on our training settings, the training time for both Qwen-VL-Chat-7B and DeepSeek-VL-Chat-7B takes around 10 hours on 1 A100. Besides, the API call costs are as follows:
  1. 1.
    Training data construction: 339,712 calls to Qwen-Max (for scoring).
  2. 2.
    Retrieve information from the web: $\sim$3000 calls to Bing/Google (for test set evaluation).
  3. 3.
    Test set inference: $\sim$3,000 calls to GPT-4o and $\sim$3,000 calls to Qwen-VL-Max
  4. 4.
    Test set LLM (Qwen-Max) evaluation: $\sim$6,000 calls (No RAG & All RAG settings)
A.5 LLM Evaluation Consistency
We report the LLM evaluation consistency on the Mix dataset with No RAG and All RAG settings in Table 9. We note that the scores of all other settings can be derived from these two settings with the "Use RAG or not" decisions. It can be observed that, under these settings, the maximum difference in the average scores between GPT-4o and Qwen-Max is within 3.
Scoring LLM Qwen-Max GPT-4o Qwen-Max GPT-4o
Setting No RAG All RAG
Ds.-VL-Chat 34.96 32.88 45.18 45.70
Qwen-VL-Chat 34.44 32.62 38.60 38.34
Qwen-VL-Max 46.54 45.88 49.26 48.45
GPT-4o 51.44 51.72 52.90 50.58
Table 9: LLM evaluation consistency between Qwen-Max and GPT-4o on Mix dataset. Scores range from 0 to 100.
A.6 Main Results on DeepSeek
@@
A.7 Supplementary Results of “Surrogate Boundary” Experiments
We provide the supplementary experimental results for Sec 4.1 where the token accuracy metrics are shown in Table 11. It can be concluded that similar conclusions can be drawn as in Sec 4.1. The experiment where DeepSeek-VL-7B-Chat is trained for surrogate boundary prediction is shown in Table 12 and 13.
@@ A.7 Supplementary Results of “Surrogate Boundary” Experiments
GPT-4o 34.52 35.96 35.17 13.50% 35.77 25.00% 35.86 54.83% 36.24 74.50%
Table 13: Knowledge Boundary model (DeepSeek-VL-7B-Chat) as a surrogate boundary identifier for other VLLMs. Results were evaluated by token accuracy.
A.8 Supplementary Results on MMMU Dataset
@@
The results in Table 14 show that the Knowledge Boundary model trained by human-labeled data helps achieve the best performance. It verifies that the aforementioned Human-labeled training data is effective. In addition, we show that our methods also exhibit substantial potential within this setting, in which both the $HKB$ and $SKB$ models predict a high search ratio over MMMU. We contend that the suboptimal performance of this dataset arises because it lies beyond the knowledge boundaries, which are challenging to validate using RAG, as delineated by the white dashed lines in Fig. 1. We present the performance of each of the 30 subjects in the MMMU validation set in Fig 4. The first row shows the LLM evaluation results, and the second shows the token accuracy metric. We can see that in most subjects “Human” setting succeeds in obtaining a higher performance than both “All RAG” and “No RAG” settings.
@@
Figure 4: Qwen-VL-Max and Qwen-VL-2 performance on the MMMU validation set with the Knowledge Boundary model trained on Human-labeled data.